OpenStudy (anonymous):

Show how to solve seven-twelfths times six-elevenths with and without cancellation.

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

\(\sf \frac{7}{12} \times \frac{6}{11}\)

OpenStudy (mathstudent55):